OpenStudy (anonymous):

Solve 4y - y + 11 = 38. a. 9 b.-9 c.3.8 d.8.6

OpenStudy (anonymous):

First you need to combine the like terms of 4y-y+11=38 and subtract 4y by y making it 3y. So then the equation will look like this: 3y+11=38 Then you want to subtract 11 from both sides 3y/3 = 27/3 Then you divide 27 / 3 Y=9